# OmniDocBench: Benchmarking Diverse PDF Document Parsing with Comprehensive Annotations

### v1

@@ 5.3 Single Algorithm Evaluation Results
Layout Detection Results. Layout detection is the first step in document parsing using pipeline tools. A robust layout detection algorithm should perform well across a variety of document types. Table 6 presents an evaluation of leading layout detection models. The DocLayout-YOLO method, which is pre-trained on diverse synthetic document data, significantly outperforms other approaches. This superiority is a key factor in MinerU’s integration of DocLayout-YOLO, contributing to its outstanding overall performance. The table also reveals that, aside from DocLayout-YOLO, other methods perform well on books and academic literature but are less effective on other document types, primarily due to a lack of pre-training on diverse documents.
@@
Table Recognition Results. Table recognition results evaluated by Tree-Edit-Distance-based Similarity (TEDS) metric are presented in Table 7. We evaluate table recognition models across three dimensions on our OmniDocBench table subset: language diversity, table frame types, and special situations. Among all models, OCR-based models demonstrate superior overall performance, with RapidTable achieving the highest scores in language diversity and maintaining stable performance across different frame types. Expert VLMs show competitive results in specific scenarios, with StructEqTable [48] excelling in no frame tables and showing better rotation robustness. General VLMs (Qwen2-VL-7B and InternVL2-8B) exhibit relatively lower but consistent performance, suggesting that while general-purpose VLMs have made progress in table understanding, they still lag behind specialized solutions.
Text Recognition Results. In the traditional OCR task, Table 8 shows that PaddleOCR leads the field, surpassing other models significantly, with GOT also performing relatively well. Selecting these two methods for the OCR module is a prudent choice.
@@
Formula Recognition Results. For formula recognition, the CDM metric provides a clear comparison in Table 9. GPT-4o, Mathpix, and UniMERNet achieve results of 86.8%, 86.6%, and 85.0%, respectively. Notably, GPT-4o excels with a recall rate of 65.5% under strict conditions requiring perfect character accuracy. Although Mathpix shows high character-level precision, it occasionally omits punctuation, such as commas, leading to a lower overall correctness rate. Nonetheless, all three models are strong candidates for formula recognition tasks.
@@ I More End-to-End Evaluation Results
Table S1 presents the evaluation results of End2End Tables grouped by Table Attributes. As it shows, most of the models perform better in English Tables rather than Chinese ones. Most models perform relatively poorly with Full Frame and No Frame tables. The accuracy of most models is affected by special conditions. Merged cells and formulas mainly test the breadth of data the model can recognize, while colored backgrounds and table rotation test their robustness. The results show that table rotation significantly impacts the accuracy of all models. Pipeline Tools perform well on more challenging tables, but colored backgrounds can affect recognition accuracy. Several Vision Language Models (VLMs) tend to perform worse on tables with merged cells, but colored backgrounds do not significantly impact table recognition accuracy.
Table S2 shows the evaluation results of End2End Text blocks grouped by Text Attributes. Almost all models have lower recognition accuracy in Chinese compared to English. Some models, such as MinerU and Marker, experience a further decrease in accuracy when recognizing mixed Chinese and English content. Complex background colors significantly affect the recognition accuracy of pipeline tools, but they have little impact on VLMs.
@@
Pipeline Tools MinerU 75.7 59.9 79.6 60.0 72.8 70.1 60.4 64.1/66.0 66.7/65.0 59.8/68.1 2.9/66.4
Marker 52.5 43.0 44.2 41.8 55.3 47.1 52.4 43.8/47.0 42.9/46.6 44.3/46.7 6.3/46.6
Mathpix 76.1 64.3 71.9 68.3 79.3 67.0 25.8 71.2/66.4 69.8/67.6 60.5/71.8 20.7/68.8
Expert Vision Models GOT-OCR 51.9 47.0 49.4 46.2 49.3 51.6 47.2 46.5/49.7 46.4/49.1 40.2/52.7 0.0/49.4
Nougat 36.5 0.4 0.0 6.3 3.6 22.2 0.0 15.1/9.1 21.2/8.9 2.8/15.3 0.0/11.4
Vision Language Models GPT4o 71.8 58.8 57.9 63.3 69.5 61.9 31.8 57.5/65.5 61.6/62.9 62.0/63.0 14.5/63.5
Qwen2-VL 57.4 62.9 72.7 70.7 64.1 48.3 57.6 49.4/68.2 48.5/64.7 63.5/60.7 41.6/61.9
InterVL2 61.5 59.3 65.9 59.7 66.5 58.7 56.2 49.6/65.9 54.4/61.6 59.4/60.6 7.3/61.1
@@
Pipeline Tools MinerU 0.123 0.206 0.742 0.163 0.147 0.513
Marker 0.267 0.389 0.499 0.339 0.389 0.497
Mathpix 0.173 0.774 0.538 0.675 0.554 0.570
@@
II Dataset Statistics and Visualization
OmniDocBench contains 981 pages, including 9 types of PDF pages, 4 types of layouts, and 3 types of languages. Some pages also include special conditions, such as watermarks. Table S3 and Figure S1 show the number of pages with each page attribute. Figures S3, S4, S5 and S6 are examples of PDF pages with different PDF types, Layout Types, and Special Issues.
Tables are also annotated with Table Attributes. There are 6 types of Table Attributes that might influent the Table Recognition accuracy: Language, Table Frame Type, Merge Cell, Colorful Background, Contain Formula, and Rotation. Table S5 shows the numbers of annotations with specific table attributes. Figures S7 and S8 are the examples of Tables with different Frames and Special Issues.
III Model Results Visualization
Figures S10, S11, S12, S13, S14, S9, S15, S16 and S17 show the examples of Good model outputs and Bad model outputs of Document Parsing among different PDF types. As it shown, different models exhibit varying performance across different PDF types. For example, MinerU detects all handwritten notes as figures, resulting in very low recognition accuracy in Notes. Marker and InternVL2 experience missed detections, leading to lower scores. InternVL2 and Qwen2-VL, in specific PDF types (such as slides or financial reports), tend to merge multi-column text.
Figures S20, S18 and S19 show the examples of Good model outputs and Bad model outputs under special issues of the PDF pages. It shows that Marker tends to generate typos when the PDF pages are fuzzy scanned or with watermarks, while GOT-OCR fails to recognize content on pages with colored backgrounds. MinerU performs well under special situations, while Mathpix occasionally generates typos.
Figures S21, S22, S23 and S24 show examples of Good model outputs and Bad model outputs for PDF pages with different layouts. MinerU has a low reading order score for single-column layouts primarily because most notes are single-column, and MinerU performs poorly in recognizing Notes, leading to a low reading order score accordingly. InternVL2 scores high in Single-Column layouts but scores poorly on Double-Column and Three-Column layouts. It is mainly due to frequent missed content recognition and errors in reading order judgment in multi-column layouts pages. MinerU’s reading order and recognition accuracy decrease with complex layouts, primarily because it incorrectly merges multiple columns during recognition.
Figures S27 and S28 show the model’s recognition ability under special issues of text. In text recognition with complex background colors, Marker may produce errors or miss content, whereas Qwen2-VL still performs well. Most models fail to recognize text when it is rotated 270 degrees. Some vision language models generate hallucinated information based on the content they can recognize.
Figures S29, S30, S31 and S32 show the examples of good and bad model results for tables with different attributes. For three-line tables, RapidTable demonstrates a good performance with accurate structure recognition, while PaddleOCR shows limitations by missing the last column in its outputs. Interestingly, in tables without frames, PaddleOCR performs well with accurate table predictions, while Qwen2-VL-7B exhibits errors in the last two columns. This indicates that the presence or absence of table frames can significantly impact different models’ performance in different ways. Rotated tables prove to be particularly challenging, with most models, including GOT-OCR, failing to recognize the table structure. However, StructEqTable shows promising results by correctly identifying most of the table content, though with a few detail errors. For tables containing formula, Qwen2-VL-7B shows more accurate table structure recognition compared to InternVL2-8B.
IV Model Settings
For pipeline tools such as MinerU, Marker, and Mathpix, default settings are used for evaluation. Specifically, MinerU with Version 0.9.3555https://github.com/opendatalab/MinerU/releases/tag/magic_pdf-0.9.3-released is employed. For Marker, Version 0.2.17666https://github.com/VikParuchuri/marker/releases/tag/v0.2.17 is evaluated. For Nougat, we utilize its 0.1.0-base model (350M). For GOT-OCR, we employ its format OCR mode to output structured data. For general VLMs, we used the GPT4o, Qwen2-VL-72B, and InternVL2-Llama3-76B by setting the do_sample===False to ensure the reproducibility.

### v2

@@ 5.2 Single Task Evaluation Results
Layout Detection Results. Layout detection is the first step in document parsing using pipeline tools. A robust layout detection algorithm should perform well across a variety of document types. Table 6 presents an evaluation of leading layout detection models. The DocLayout-YOLO method, which is pre-trained on diverse synthetic document data, significantly outperforms other approaches. This superiority is a key factor in MinerU’s integration of DocLayout-YOLO, contributing to its outstanding overall performance. Other methods perform well on books and academic literature but struggle with more diverse formats due to limited training data.
@@
Table Recognition Results. In Table 7, We evaluate table recognition models across three dimensions on our OmniDocBench table subset: language diversity, table frame types, and special situations. Among all models, OCR-based models demonstrate superior overall performance, with RapidTable achieving the highest scores in language diversity and maintaining stable performance across different frame types. Expert VLMs show competitive results in specific scenarios, with StructEqTable [55] excelling in no-frame tables and showing better rotation robustness. General VLMs (Qwen2-VL-7B and InternVL2-8B) exhibit relatively lower but consistent performance, suggesting that while general-purpose VLMs have made progress in table understanding, they still lag behind specialized solutions.
Text Recognition Results. Table 8 compares OCR tools across languages, backgrounds, and rotations using Edit Distance. PaddleOCR outperforms all competitors, followed by GOT-OCR and Mathpix. General VLMs struggle to handle text rotation or mixed-language scenarios.
@@
Formula Recognition Results. Table 9 presents results on formula parsing, using CDM, BLEU, and normalized Edit Distance. GPT-4o, Mathpix, and UniMERNet achieve results of 86.8%, 86.6%, and 85.0%, respectively. Notably, GPT-4o excels with a recall rate of 65.5% under strict conditions requiring perfect character accuracy. Although Mathpix shows high character-level precision, it occasionally omits punctuation, such as commas, leading to a lower overall correctness rate. Nonetheless, all three models are strong candidates for formula recognition tasks.
@@ I More End-to-End Evaluation Results
Table S1 presents the evaluation results of End2End Tables grouped by Table Attributes. As it shows, most of the models perform better in English Tables rather than Chinese ones. Most models perform relatively poorly with Full Frame and No Frame tables. The accuracy of most models is affected by special conditions. Merged cells and formulas mainly test the breadth of data the model can recognize, while colored backgrounds and table rotation test their robustness. The results show that table rotation significantly impacts the accuracy of all models. Pipeline Tools’ performance would not be affected by more challenging tables (e.g., merge cell), but colored backgrounds can affect recognition accuracy. Several Vision Language Models (VLMs) tend to perform worse on tables with merged cells, but colored backgrounds do not significantly impact table recognition accuracy.
Table S2 shows the evaluation results of End2End Text blocks grouped by Text Attributes. Almost all models have lower recognition accuracy in Chinese compared to English. Some models, such as MinerU and Marker, experience a further decrease in accuracy when recognizing mixed Chinese and English content. The main reason is that minerU’s text recognition module is PaddleOCR model. According to the performance of the PaddleOCR model in text recognition module, its accuracy will decline in the case of mixed language. Moreover, complex background colors significantly affect the recognition accuracy of pipeline tools, but it has only little impact on accuracy for VLMs.
@@
Pipeline Tools MinerU 75.1 59.3 79.1 59.4 71.6 69.7 60.0 63.6/65.3 66.0/64.4 59.2/67.5 3.0/65.8
Marker 64.9 47.3 49.8 44.5 61.8 59.0 63.6 52.6/52.7 53.2/52.5 48.0/54.9 35.5/52.9
Mathpix 75.4 63.2 71.3 67.4 77.3 66.3 25.5 70.3/65.4 68.7/66.7 59.7/70.8 19.2/67.9
Expert Vision Models GOT-OCR 51.7 46.2 49.0 45.5 48.3 51.3 46.2 46.0/48.9 45.7/48.4 39.8/51.9 0.0/48.7
Nougat 36.2 0.3 0.0 6.1 3.5 22.1 0.0 15.0/8.9 21/8.7 2.6/15.2 0.0/11.2
Vision Language Models GPT4o 71.1 58.0 57.3 62.5 68.7 61.3 31.2 56.8/64.7 60.8/62.2 61.4/62.2 14.2/62.7
Qwen2-VL-72B 73.2 75.1 76.1 72.0 79.0 77.5 63.2 67.9/78.1 71.6/75.3 77.9/72.9 42.7/75.1
InterVL2-76B 60.9 58.5 65.4 58.8 65.3 58.3 55.6 49.0/65.1 53.3/60.9 58.8/59.8 6.9/60.3
@@
Pipeline Tools MinerU 0.124 0.234 0.742 0.188 0.15 0.514
Marker 0.163 0.379 0.747 0.303 0.396 0.594
Mathpix 0.175 0.793 0.538 0.698 0.587 0.583
@@
II Dataset Statistics and Visualization
OmniDocBench contains 981 pages, including 9 types of PDF pages, 4 types of layouts, 3 types of languages, and 3 special issues in visual degradations (e.g., watermarks). Table S3 and Figure S1 show the number of pages with each page attribute. Figures S5, S6, S7 and S8 are examples of PDF pages with different PDF types, Layout Types, and Special Issues.
Tables are also annotated with Table Attributes. There are 6 types of Table Attributes that might influent the Table Recognition accuracy: Language, Table Frame Type, Merge Cell, Colorful Background, Contain Formula, and Rotation. Table S5 shows the numbers of annotations with specific table attributes. Figures S9 and S10 are the examples of Tables with different Frames and Special Issues.
III Discussion on Model Predictions
Conclusion Combining scattered results from tasks and sub-attributes, it can be concluded that pipeline tools and expert models have better performance on common data like academic papers and challenging cases such as tables with merged cells compared to VLMs. However, VLMs demonstrate stronger generalization on uncommon PDF types like slides and exam papers, and they show greater robustness in special page situations, such as fuzzy scans. The low accuracy of VLMs is mainly due to:1) Missing Content in dense pages(Figure S11); 2) Hallucinations in hard-to-recognize pages(Figure S30). The low accuracy of Pipeline tools mainly due to: 1) Lower robustness in special page situations, e.g., watermark(Figure S21); 2) Weak generalization on uncommon PDF types, e.g., handwriting notes(Figure S16).
Figures S12, S13, S14, S15, S16, S11, S17, S18 and S19 show the examples of Good model outputs and Bad model outputs of Document Parsing among different PDF types. As it shown, different models exhibit varying performance across different PDF types. For example, MinerU detects all handwritten notes as figures, resulting in very low recognition accuracy in Notes. Marker and InternVL2 experience missed detections, leading to lower scores. InternVL2 and Qwen2-VL, in specific PDF types (such as slides or financial reports), tend to merge multi-column text.
Figures S22, S20 and S21 show the examples of Good model outputs and Bad model outputs under special issues of the PDF pages. It shows that Marker tends to generate typos when the PDF pages are fuzzy scanned or with watermarks, while GOT-OCR fails to recognize content on pages with colored backgrounds. MinerU performs well under special situations, while Mathpix occasionally generates typos.
Figures S23, S24, S25 and S26 show examples of Good model outputs and Bad model outputs for PDF pages with different layouts. MinerU has a low reading order score for single-column layouts primarily because most notes are single-column, and MinerU performs poorly in recognizing Notes, leading to a low reading order score accordingly. InternVL2 scores high in Single-Column layouts but scores poorly on Double-Column and Three-Column layouts. It is mainly due to frequent missed content recognition and errors in reading order judgment in multi-column layouts pages. MinerU’s reading order and recognition accuracy decrease with complex layouts, primarily because it incorrectly merges multiple columns during recognition.
Figures S29 and S30 show the model’s recognition ability under special issues of text. In text recognition with complex background colors, Marker may produce errors or miss content, whereas Qwen2-VL still performs well. Most models fail to recognize text when it is rotated 270 degrees. Some vision language models generate hallucinated information based on the content they can recognize.
Figures S31, S32, S33 and S34 show the examples of good and bad model results for tables with different attributes. For three-line tables, RapidTable demonstrates a good performance with accurate structure recognition, while PaddleOCR shows limitations by missing the last column in its outputs. Interestingly, in tables without frames, PaddleOCR performs well with accurate table predictions, while Qwen2-VL-7B exhibits errors in the last two columns. This indicates that the presence or absence of table frames can significantly impact different models’ performance in different ways. Rotated tables prove to be particularly challenging, with most models, including GOT-OCR, failing to recognize the table structure. However, StructEqTable shows promising results by correctly identifying most of the table content, though with a few detail errors. For tables containing formula, Qwen2-VL-7B shows more accurate table structure recognition compared to InternVL2-8B.
IV Model Settings
For pipeline tools such as MinerU, Marker, and Mathpix, default settings are used for evaluation. Specifically, MinerU with Version 0.9.3888https://github.com/opendatalab/MinerU/releases/tag/magic_pdf-0.9.3-released is employed. For Marker, Version 1.2.3999https://github.com/VikParuchuri/marker/releases/tag/v1.2.3 is evaluated. For Nougat, we utilize its 0.1.0-base model (350M). For GOT-OCR, we employ its format OCR mode to output structured data.
For general VLMs, we used the GPT4o, Qwen2-VL-72B, and InternVL2-Llama3-76B by setting the do_sample===False to ensure the reproducibility. After testing the different setting of max_token, the best setting is chosen for each VLMs. Specifically, max_token===32000 is set for Qwen2-VL-72B, and max_token===4096 is set for InternVL2-Llama3-76B. For GPT-4o, the default setting is used.
V More Details on Methods
Ignore handling. The purpose of this process is to avoid fluctuations in accuracy caused by the lack of uniformity in the output standards among document parsing algorithm. (1) Some algorithm (e.g., GPT-OCR, Qwen2-VL) tends to remove headers and footers, while others (e.g., GPT4o) prefers to retain them ( Figure S3). (2) Moreover, the reading order mismatch cause by captions and footnotes is also considered. For example, Nougat would put the image captions in the end of the page content( Figure S4), while others tend to put the image captions in human reading order.
Ignore handling is to minimize the impact of varying standards of document parsing on evaluation. Our evaluation dataset aims to more fairly assess the parsing accuracy of various algorithms, and these trivial issues regarding standards are not within our scope of consideration.
@@ V More Details on Methods
Figure S3: The Vary Standards in parsing Header, Footers, and so on.
Figure S4: The Vary Standards in parsing Captions.